OpenStudy (anonymous):

(square root of x) x − (square root of "x-7") x − 7= 1

OpenStudy (anonymous):

its pretty much the square root of x minus the square root of x minus 7 equals 1. please solve

OpenStudy (anonymous):

\[\sqrt x -\sqrt {x-7}=1\]

OpenStudy (anonymous):

right?

OpenStudy (aum):

Is the equation: \(\sqrt x -\sqrt {x-7}=1\) ?

OpenStudy (anonymous):

it is 16 i guess and checked.

OpenStudy (aum):

But if you answer if that equation is correct or not you can learn the proper way to solve this problem without having to guess.